# Rethinking Graph Structure Learning in the Era of LLMs

### v1

@@ Abstract
Recently, the emergence of large language models (LLMs) has prompted researchers to explore the integration of language descriptions into graphs, aiming to enhance model encoding capabilities from a data-centric perspective. This graph representation is called text-attributed graphs (TAGs). A review of prior advancements highlights that graph structure learning (GSL) is a pivotal technique for improving data utility, making it highly relevant to efficient TAG learning. However, most GSL methods are tailored for traditional graphs without textual information, underscoring the necessity of developing a new GSL paradigm.
Despite clear motivations, it remains challenging: (1) How can we define a reasonable optimization objective for GSL in the era of LLMs, considering the massive parameters in LLM? (2) How can we design an efficient model architecture that enables seamless integration of LLM for this optimization objective? For Question 1, we reformulate existing GSL optimization objectives as a tree optimization framework, shifting the focus from obtaining a well-trained edge predictor to a language-aware tree sampler. For Question 2, we propose decoupled and training-free model design principles for LLM integration, shifting the focus from computation-intensive fine-tuning to more efficient inference. Based on this, we propose Large Language and Tree Assistant (LLaTA), which leverages tree-based LLM in-context learning to enhance the understanding of topology and text, enabling reliable inference and generating improved graph structure. Extensive experiments on 10 TAG datasets demonstrate that LLaTA enjoys flexibility - incorporated with any backbone; scalability - outperforms other LLM-based GSL methods in terms of running efficiency; effectiveness - achieves SOTA performance.
Graph Structure Learning, Large Language Model, ICML
@@ 1 Introduction
In recent years, with the prevalence of large language models (LLMs) in graph ML (Yan et al., 2023; Chen et al., 2024), text-attributed graphs (TAGs) have emerged as a novel graph representation. This data structure leverages textual information to provide fine-grained descriptions of graphs, bringing systemic transformation for graph learning (Li et al., 2024). Inspired by this, achieving efficient and robust learning of TAGs has become an urgent priority.
To achieve this, graph structure learning (GSL) is a promising approach that enhances data utility for vanilla graph neural networks (GNNs). However, most prevalent GSL methods (Zhiyao et al., 2024) are designed for traditional graphs and fail to process rich textual information, leading to sub-optimal performance. Despite significant efforts by recent LLM-based GSL methods, such as GraphEdit (Guo et al., 2024) and LLM4RGNN (Zhang et al., 2024), their optimization objectives and model architectures still adhere to traditional GSL paradigms, inheriting unnecessary complexities. Specifically, these prominent GSL approaches (Zhao et al., 2021; Wu et al., 2023; Jiang et al., 2024; Guo et al., 2024; Zhang et al., 2024) rely on carefully designed loss functions to dynamically maintain a graph learner (i.e., structure optimizer) coupled with well-defined instruction datasets or a specific GNN backbone (i.e., the downstream encoder with improved structure). Based on this, we have the following 2 observations and seek to explore them to gain key insights into developing a new GSL paradigm in the era of LLMs.
Observation 1 (Optimization Objective): The graph learner often serves as an edge predictor, heavily relying on end-to-end training with a specific downstream backbone. In the era of LLMs, LLMs will play a crucial role in graph learners. In this context, performing full-parameter LLM training to obtain an edge predictor is infeasible. While two recent parameter-efficient fine-tuning methods have been proposed for GSL, the complexity of constructing instruction datasets and tuning significantly hinders their efficiency. This leads to Question 1: How can we define a reasonable optimization objective for GSL in the era of LLMs?
@@
Observation 2 (Model Architecture): To obtain a well-trained edge predictor, the existing graph learner is often coupled with customized instruction datasets or a specific backbone, with gradient supervision for model parameter updates in a collaborative manner. The complexity of this model architecture hinders the adaptability of the improved graph structure to real-world scenarios (i.e., deployment efficiency, instruction-free, and backbone-free). In the era of LLMs, LLMs possess remarkable in-context learning capability with textual information, paving the way for efficiently obtaining improved structure. Based on this and Obs.1, we adopt a tree-oriented in-context learning approach. However, we must address Question 2: How can LLM be seamlessly integrated into the model architecture for efficient GSL?
Answer 2: In Sec. 3.2, we establish the decoupled and training-free model design principles by revisiting existing GSL, emphasizing reliable LLM inference over fine-tuning.
@@
Our contributions. (1) New Perspective. We systematically review existing GSL methods and provide empirical investigations, revealing the challenges and opportunities for GSL in the era of LLMs. (2) Innovative Approach. We reformulate the existing GSL and propose a tree-based optimization framework with model design principles. Based on this, we introduce LLaTA, which seamlessly integrates both topology and text insights by tree-driven prompts to facilitate LLM in-context learning and generate improved graph structure. (3) SOTA Performance. Extensive experiments demonstrate the superior performance of LLaTA. It outperforms recent LLM-based GraphEdit (Guo et al., 2024) by 1.3%-2.5% in accuracy while running 3.0h-4.0h faster.
@@ 2.2 Complexity Metrics of Graph Topology
Structural Entropy. Motivated by Shannon entropy (Shannon, 1948), structural entropy (SE) (Li & Pan, 2016) is an effective measurement for quantifying the dynamic complexity of graph topology. By minimizing SE, we can reduce structure uncertainty and capture inherent patterns, thereby supporting downstream tasks in a robust and interpretable manner. In other words, this approach uncovers meaningful structural patterns and ensures that it aligns more effectively with task-specific requirements. This has made it a pivotal tool for GNNs, gaining significant attention in recent years.
@@ 2.3 Related Works
SE-based Methods. Existing SE-based methods leverage the well-defined structural encoding tree to perform a wide range of graph-based downstream tasks such as node clustering (Pan et al., 2021), community detection (Liu et al., 2019), multi-layer coarsening (Wu et al., 2022), and embedding dimension estimation (Yang et al., 2023). Recent SEGSL (Zou et al., 2023) enhances edge connectivity and quality using this structural encoding tree, but it relies on end-to-end tree-based training without incorporating LLMs.
Traditional GSL. These methods can be categorized into metric-based (Chen et al., 2020; Li et al., 2022; Wu et al., 2023), probabilistic sampling (Zhao et al., 2021; Luo et al., 2021; Liu et al., 2022a), and directly learnable methods (Jin et al., 2020; Liu et al., 2022b). Specifically, they refine graph structure using well-designed measurement (e.g., homophily and connectivity), edge re-weight, random sampling, or direct topology optimization without textual information.
LLM-based GSL. GraphEdit (Guo et al., 2024) improves node connectivity by adding neighbors through a well-trained edge predictor and refines the improved graph structure by evaluating node relevance using a fine-tuned LLM. LLM4RGNN (Zhang et al., 2024) leverages a fine-tuned LLM to infer node relevance, which is used to train an edge predictor. Compared to improving the graph structure, it focuses more on robustness against graph adversarial attacks.
@@ 3.1 GSL Optimization Objectives Reformulation
As highlighted by Obs.1 in Sec. 1, the optimization objective of existing GSL is to obtain a well-trained edge predictor as the graph learner. This requires coupling the graph learner with a specific backbone, relying on training in a collaborative manner. Despite its effectiveness, this optimization objective becomes inefficient when applied to LLMs due to their massive trainable parameters. During our investigation, we noted that existing LLM-based GSL methods adopt instruction fine-tuning to address this full-parameter training issue. However, fully understanding intricate graph topology and constructing tailored instruction datasets often demand considerable effort, as evidenced by their complex workflows (e.g., GraphEdit and LLM4RGNN). Therefore, we reformulate GSL and propose a new tree-based optimization framework rather than relying on an edge predictor, as illustrated in Fig. 1 (a)-(c). The core of our framework is to obtain a language-aware tree sampler for efficient GSL.
Specifically, (a) Tree Construction first reorganizes the original graph to assist LLMs in understanding the topology. Based on this, (b) Tree Optimization leverages textual information and LLM to perform tree optimization. Finally, (c) Improved Structure is generated by a language-aware tree sampler. Please refer to Sec.3.2 and Sec.4 for details on the instantiation of the above framework to propose LLaTA.
@@ 3.2 Model Architecture Review
Coupled Paradigm. The graph learner and backbone are tightly coupled (Chen et al., 2020; Fatemi et al., 2021; Wu et al., 2023) to learn improved structure by specific downstream tasks (e.g., node, link, and graph level). Limitations: (1) Unstable Performance: This task- and backbone-specific model architecture may limit the generalizability of the improved structure and raise fairness concerns due to the potential data quality issues. (2) Inflexible Deployment: Switching the downstream backbone or task requires retraining, which significantly limits its adaptability and scalability.
Decoupled Paradigm. The graph learner and downstream backbone are trained independently (Liu et al., 2022b, a; Bi et al., 2024), avoiding co-training, thereby enabling better compatibility with LLMs for GSL. We observe that recent LLM-based GSL methods also tend to adopt this decoupled paradigm (Guo et al., 2024; Zhang et al., 2024). However, these methods exhibit a different form of coupling (i.e., fine-tuning based on the pre-defined instruction datasets), which leads to the following issues: (1) They fail to directly and jointly incorporate topology and text, relying instead on instruction datasets. Moreover, they are highly sensitive to the quality of these datasets. (2) They rely on complex mechanisms and significant time to construct instruction datasets and fine-tuning. Therefore, from a model design perspective, we recommend adopting a comprehensive decoupled paradigm to integrate GSL and LLMs, minimizing reliance on fine-tuning and instead prioritizing more efficient and reliable inference with high-quality in-context prompts.
To achieve this, we emphasize leveraging the inherent in-context learning capability of LLMs to enable a comprehensive understanding of topology and text for reliable inference. Specifically, (a) Tree Construction: we first construct topology-aware in-context based on the structural encoding tree, which serves as high-quality prompts to enhance the LLM’s understanding of graph topology. Based on these tree-driven prompts, (b) Tree Optimization: we leverage the language descriptions of nodes to capture their semantic relationships through LLM inference, which is then utilized to reallocate leaf dependencies for optimizing the tree structure. The hierarchical structure of this optimized tree reveals topology and text insights from the original graph, providing a solid foundation for GSL. Finally, (c) Improved Structure: we perform leaf-oriented two-step sampling to identify the current node and a candidate node set for edge addition or removal, resulting in improved graph structure for any downstream scenario. For deeper insights into the leaf-oriented sampler within GSL, please refer to Appendix B.
@@ 3.3 Empirical Investigation
To demonstrate the effectiveness of our proposal in Sec. 3.1-3.2, we present empirical results in Fig. 1(d). Due to space constraints, detailed experimental settings and analysis are provided in Appendix C. In our reports, over-smoothing ($\downarrow$) reflects the distinguishability of node embeddings, while over-squashing ($\uparrow$) indicates the ability of nodes to perceive distant ones. These metrics are widely used in recent GSL studies to quantify the quality of improved structure, as they not only directly impact downstream accuracy ($\uparrow$) but also evaluate method robustness. They demonstrate that LLM-based GSL methods generate higher-quality structures than traditional ones, underscoring the necessity of LLMs. Furthermore, LLaTA achieves the best performance, validating the effectiveness of our proposed new GSL paradigm.
Figure 2: (Left) The overview of LLaTA; (Right) The detailed pipeline of LLaTA. LLaTA achieves training-free GSL through language-aware tree sampler, which is defined by topology-aware tree prompts and reliable LLM inference.
@@ 4.1 Overview of LLaTA
We instantiate the three components outlined in Fig. 1 and propose LLaTA, shown in Fig. 2. The correspondences are: (a)-(1) Topology-aware In-context Construction: We leverage a structural encoding tree to reorganize the graph and utilize low-level tree communities to construct topology-aware prompts. (b)-(2) Tree-prompted LLM Inference: We enable efficient LLM in-context learning with these prompts and node texts, allowing the model to fully capture both topology and text insights during inference. The inference results are then used to optimize the tree through leaf allocation. (c)-(3) Leaf-oriented Two-step Sampling: We utilize the language-aware tree sampler to facilitate edge addition and removal and obtain improved graph structure.
@@ 4.2 Topology-aware In-context Construction
Motivation. To address the potential performance decline caused by the absence of fine-tuning and obtain reliable inference, we enable efficient in-context learning with high-quality prompts. Based on Sec. 2.2, structural entropy is a promising strategy for constructing these prompts, as it offers an interpretable theory for capturing topology insights.
To establish a hierarchical encoding tree that simulates the natural evolution of graph topology, we constrain the height of this tree to $K$ and minimize $K$-dimensional SE proposed by (Li & Pan, 2016). The optimization objective definition:
$\mathcal{T}^{\star}\!\!=\!\arg\!\min_{\mathcal{T}}\mathcal{H}^{\mathcal{T}}(\mathcal{G})\equiv-\!\!\!\!\!\!\!\sum_{\forall t\in\mathcal{T},t\neq\lambda}\!\!\!\frac{g_{t}}{\operatorname{vol}(\mathcal{V})}\log\!\frac{\operatorname{vol}\left(t\right)}{\operatorname{vol}\left(t^{+}\right)},$ (1)
where $\operatorname{vol}(\mathcal{G})$ is the sum of the degrees of all nodes in $\mathcal{G}$, $t^{+}$ is the parent of $t$, $\lambda$ is the root node, $\mathcal{T}^{\star}$ denotes the optimized tree that minimizes its structural entropy $\mathcal{H}^{\mathcal{T}}(\mathcal{G})$. The two meta-operations for tree construction are as follows.
Node Combining merges sibling nodes into a new parent node, replacing their original parent in the encoding tree.
Node Lifting raises a node to the same level as its parent node, removing its original parent if it becomes childless.
Based on this, the core of our employed greedy algorithm is as follows: (1) Initialization. The structural encoding tree is initialized with 1 height. Then, a root node is created and each original graph node is a leaf node. (2) Node Combination. Pairs of nodes in the tree are iteratively combined to minimize $\mathcal{H}^{\mathcal{T}}(\mathcal{G})$ at each step. This process continues until the root node has at most two children. (3) Node Lifting. Nodes are iteratively lifted to minimize $\mathcal{H}^{\mathcal{T}}(\mathcal{G})$ after each operation. This process is repeated until the tree height is reduced to $K$, resulting in a final encoding tree $\mathcal{T}^{\star}$ with height $K$. Due to space constraints, the complete algorithm description is provided in Appendix D.1.

### v2

@@ Abstract
Recently, the emergence of LLMs has prompted researchers to integrate language descriptions into graphs, aiming to enhance model encoding capabilities from a data-centric perspective. This graph representation is called text-attributed graphs (TAGs). A review of prior advancements highlights that graph structure learning (GSL) is a pivotal technique for improving data utility, making it highly relevant to efficient TAG learning. However, most GSL methods are tailored for traditional graphs without textual information, underscoring the necessity of developing a new GSL paradigm. Despite clear motivations, it remains challenging: (1) How can we define a reasonable optimization objective for GSL in the era of LLMs, considering the massive parameters in LLM? (2) How can we design an efficient model architecture that enables seamless integration of LLM for this optimization objective? For Question 1, we reformulate existing GSL optimization objectives as a tree optimization framework, shifting the focus from obtaining a well-trained edge predictor to a language-aware tree sampler. For Question 2, we propose decoupled and training-free model design principles for LLM integration, shifting the focus from computation-intensive fine-tuning to more efficient inference. Based on this, we propose Large Language and Tree Assistant (LLaTA), which leverages tree-based LLM in-context learning to enhance the understanding of topology and text, enabling reliable inference and generating improved graph structure. Extensive experiments on 10 datasets demonstrate that LLaTA enjoys flexibility-incorporated with any backbone; scalability-outperforms other LLM-enhanced graph learning methods; effectiveness-achieves SOTA predictive performance.
@@ 1 Introduction
In recent years, the rise of LLMs in graph ML [29, 3] has led to the emergence of text-attributed graphs (TAGs) as a novel data representation. This data structure leverages textual information to provide fine-grained descriptions of graphs, driving a significant shift in data-centric graph learning [15]. Inspired by this, efficiently and robustly learning TAGs has become an urgent priority.
To achieve this, graph structure learning (GSL) is a promising approach that enhances data utility for vanilla graph neural networks (GNNs). However, most existing prevalent GSL approaches [35] are designed for traditional graphs and fail to process rich textual information, leading to sub-optimal performance. Despite significant efforts by recent LLM-based GSL methods, such as GraphEdit [7] and LLM4RGNN [31], their optimization objectives and model architectures still adhere to traditional GSL paradigms, inheriting unnecessary complexities. Specifically, these prominent GSL approaches [32, 28, 10, 7, 31] rely on carefully designed loss functions to dynamically maintain a graph learner (i.e., structure optimizer) coupled with well-defined instruction datasets or a specific GNN backbone (i.e., improved structure tailored for the specific downstream encoder). Based on this, we have the following 2 observations and seek to explore them to gain key insights into developing a new GSL paradigm in the era of LLMs.
Observation 1 (Optimization Objective): The graph learner often serves as an edge predictor, heavily relying on end-to-end training with a specific downstream backbone. In the era of LLMs, LLMs will play a crucial role in graph learners. In this context, performing full-parameter LLM training to obtain an edge predictor is infeasible. While two recent fine-tuning methods have been proposed, the complexity of constructing instruction datasets and tuning still hinders their efficiency. This leads to Question 1: How can we define a reasonable optimization objective for LLM-enhanced GSL?
@@
Observation 2 (Model Architecture): To obtain a well-trained edge predictor, the existing graph learner is often coupled with customized instruction datasets or a specific graph learning backbone, with gradient supervision for model parameter updates in a collaborative manner. The complexity of this model architecture hinders the adaptability of the improved graph structure to real-world scenarios (i.e., deployment efficiency, instruction-free, and backbone-free). In the era of LLMs, LLMs possess remarkable in-context learning capability with textual information, paving the way for efficiently obtaining improved structure. Based on this and Observation 1, we adopt a tree-oriented in-context learning approach. However, we must address Question 2: How can LLM be seamlessly integrated into the model architecture for efficient GSL?
Answer 2: In Sec. 3.2, we establish the decoupled and training-free model design principles by revisiting existing GSL, emphasizing efficient LLM inference over computation-intensive fine-tuning.
@@
Our contributions. (1) New Perspective. We systematically review existing GSL methods and provide empirical investigations, revealing the challenges and opportunities for GSL in the era of LLMs. (2) Innovative Approach. We reformulate the existing GSL and propose a tree-based optimization framework with model design principles. Based on this, we introduce LLaTA, which seamlessly integrates both topology and text insights by tree-driven prompts to facilitate LLM in-context learning and generate improved graph structure, with complete theoretical support. (3) SOTA Performance. Extensive experiments demonstrate the superior performance of LLaTA. It outperforms recent LLM-based methods by 1.3%-2.5% in accuracy while running 3.0h-9.7h faster.
@@ 2.2 Complexity Metrics of Graph Topology
Structural Entropy. Motivated by Shannon entropy [23], structural entropy (SE) [13] is an effective measurement for quantifying the dynamic complexity of graph topology. By minimizing SE, we can reduce structure uncertainty and capture inherent patterns, thereby supporting downstream tasks in a robust and interpretable manner. In other words, this approach uncovers meaningful structural patterns and ensures that it aligns more effectively with task-specific requirements. This has made it a pivotal tool for GNNs, gaining significant attention in recent years.
@@ 2.3 Related Works
SE-based Methods. Existing SE-based methods leverage the well-defined structural encoding tree to perform a wide range of graph-based downstream tasks such as node clustering [22], community detection [17], multi-layer coarsening [27], and embedding dimension estimation [30]. Recent SEGSL [37] enhances edge connectivity and quality using this structural encoding tree, but it relies on end-to-end tree-based training without incorporating LLMs.
Traditional GSL. These methods can be categorized into metric-based [2, 14, 28], probabilistic sampling [32, 19, 16], and directly learnable methods [11, 18]. Specifically, they refine graph structure using well-designed measurement (e.g., homophily and connectivity), edge re-weight, random sampling, or direct topology optimization without textual information.
LLM-based GSL. GraphEdit [7] improves node connectivity by adding neighbors through a well-trained edge predictor and refines the improved graph structure by evaluating node relevance using a fine-tuned LLM. LLM4RGNN [31] leverages a fine-tuned LLM to infer node relevance, which is used to train an edge predictor. Compared to improving the graph structure, it focuses more on robustness against graph adversarial attacks.
@@ 3.1 GSL Optimization Objectives Reformulation
As highlighted by Obs.1 in Sec. 1, existing GSL methods aim to train an effective edge predictor as the graph learner. This typically requires coupling the graph learner with a specific backbone and jointly optimizing them. Although effective, this objective becomes inefficient for LLMs due to their large number of parameters. Our investigation shows that current LLM-based GSL approaches often adopt instruction fine-tuning to mitigate full-parameter training. However, understanding complex graph structures and crafting appropriate instruction datasets can be labor-intensive, as reflected in their elaborate workflows [7, 31]. To address this, we reformulate GSL and introduce a tree-based optimization framework that avoids edge predictors, as illustrated in Fig. 1 (a)-(c). The core idea is to develop a language-aware tree sampler for efficient GSL.
Specifically, (a) Tree Construction first reorganizes the original graph to assist LLMs in understanding the topology. Based on this, (b) Tree Optimization leverages textual information and LLM to perform tree optimization. Finally, (c) Improved Structure is generated by a language-aware tree sampler. Please refer to Sec.3.2 and Sec.4 for implementation of the above framework in LLaTA.
@@ 3.2 Model Architecture Review
The Coupled Paradigm tightly integrates the graph learner and backbone for task-specific learning. However, it faces limitations such as unstable performance due to restricted generalizability and fairness concerns, and inflexible deployment as switching tasks or backbones requires retraining. In contrast, the Decoupled Paradigm trains the graph learner and backbone independently, offering better compatibility with LLMs for GSL. Although recent LLM-based GSL methods adopt this approach, they still face challenges such as indirect integration of topology and text, reliance on instruction datasets, and the complexity of fine-tuning. For a more efficient and reliable approach, we recommend a fully decoupled paradigm, reducing fine-tuning reliance and focusing on high-quality in-context prompts. Detailed discussion is provided in the Appendix C.
To achieve this, we leverage the in-context learning ability of LLMs to jointly understand graph topology and node text for reliable inference. Specifically, (a) Tree Construction: we first build a structural encoding tree to generate topology-aware in-context prompts, enhancing the LLM’s understanding of graph structure. Then, (b) Tree Optimization: based on these prompts, we use LLMs to infer semantic relationships among nodes from their textual descriptions, and reassign leaf dependencies to optimize the tree structure. The resulting hierarchical tree integrates both structural and textual information, forming a strong basis for GSL. Finally, (c) Improved Structure: we apply a leaf-oriented two-step sampling to select the current node and a candidate set for edge modification, yielding an improved graph for any downstream scenario. For deeper insights into the leaf-oriented sampler within GSL, please refer to Appendix B.
@@ 3.3 Empirical Investigation
To demonstrate the effectiveness of our proposal in Sec. 3.1-3.2, we present empirical results in Fig. 1(d). Due to space constraints, detailed experimental settings and analysis are provided in Appendix D. In our reports, over-smoothing ($\downarrow$) reflects the distinguishability of node embeddings, while over-squashing ($\uparrow$) indicates the ability of nodes to perceive distant ones. These metrics are widely used in recent GSL studies to quantify the quality of improved structure, as they not only directly impact downstream accuracy ($\uparrow$) but also evaluate method robustness. They demonstrate that LLM-based GSL methods generate higher-quality structures than traditional ones, underscoring the necessity of LLMs. Furthermore, LLaTA achieves the best performance, validating the effectiveness of our proposed new GSL paradigm.
@@ 4.1 Overview of LLaTA
We instantiate the three components outlined in Fig. 1 and propose LLaTA, shown in Fig. 2. The correspondences are: (a)-(1) Topology-aware In-context, (b)-(2) Tree-prompted LLM Inference and (c)-(3) Leaf-oriented Two-step Sampling.
Figure 2: (Left) The overview of LLaTA; (Right) The detailed pipeline of LLaTA, which includes: topology-aware tree prompts, reliable LLM inference and language-aware tree sampler.
@@ 4.2 Topology-aware In-context Construction
Motivation. To address the potential performance decline caused by the absence of fine-tuning and obtain reliable inference, we enable efficient in-context learning with high-quality prompts. Based on Sec. 2.2, structural entropy is a promising strategy for constructing these prompts, as it offers an interpretable theory for capturing both local and global topology insights.
To establish a hierarchical encoding tree that simulates the natural evolution of graph topology, we constrain the height of this tree to $K$ and minimize $K$-dimensional SE proposed by [13]. The optimization objective definition:
$\mathcal{T}^{\star}\!\!=\!\arg\!\min_{\mathcal{T}}\mathcal{H}^{\mathcal{T}}(\mathcal{G}),\>\>\>\>\mathcal{H}^{\mathcal{T}}(\mathcal{G})=\!\!\!\!\!\sum_{\phi\in\mathcal{T},\phi\neq\lambda}\!\!\!\!\mathcal{H}^{\mathcal{T}}(\mathcal{G},\phi)=-\!\!\!\!\!\sum_{\phi\in\mathcal{T},\phi\neq\lambda}\!\!\!\frac{g_{\phi}}{\operatorname{vol}(\mathcal{G})}\log\!\frac{\operatorname{vol}\left(\phi\right)}{\operatorname{vol}\left(\phi^{+}\right)},$ (1)
where $\operatorname{vol}(\mathcal{G})$ is the sum of the degrees of all nodes in $\mathcal{G}$, $\operatorname{vol}(\phi)$ is the sum of the degrees of nodes in $\phi$, $\phi^{+}$ is the parent of $\phi$, $\lambda$ is the root node, $\mathcal{T}^{\star}$ denotes the optimized tree that minimizes its structural entropy $\mathcal{H}^{\mathcal{T}}(\mathcal{G})$. We employ a greedy algorithm to minimize the structural entropy $\mathcal{H}^{\mathcal{T}}(\mathcal{G})$ and construct a hierarchical encoding tree $\mathcal{T}^{\star}$ with a predefined height $K$. The algorithm starts by initializing a tree of height 1, where each graph node is a leaf under a common root. It then iteratively combines node pairs to reduce $\mathcal{H}^{\mathcal{T}}(\mathcal{G})$, followed by node lifting operations to further optimize the structure. The node lifting process continues until the tree height is reduced to $K$. The full algorithm and the operations involved is detailed in Appendix E.1.
Theorem 1 (Topological Information Capturing Properties of Encoding Tree).
Given an encoding tree $\mathcal{T}$ and a non-leaf node $\phi\in\mathcal{T}$, the error of topological information $\varepsilon^{h}(\phi)$ in community $\mathcal{C}_{\phi}$ is upper bounded by: $\frac{g_{\phi}}{2m}\log_{2}\frac{\operatorname{vol}(\phi^{+})}{g_{\phi}}$, and $\varepsilon^{h}(\phi)$ gradually decreases as the community level descends.
Theorem 2 (Implicit Global Constraints in Low-Level Communities).
In a structural encoding tree $\mathcal{T}$, each low-level community $\mathcal{C}^{\ell}$ captures localized topology but implicitly retains global structure constraints due to the hierarchical, random-walk-based formulation of structural entropy.
Drawing on Theorems 1 and  2, we infer that low-level communities in the encoding tree provide a more faithful representation of the graph’s topology, as they simultaneously preserve local connectivity and global structural constraints. Accordingly, for each leaf node $\alpha$, we assign its associated low-level community $\mathcal{C}^{\ell}(\alpha)$ as a topological context to enhance the accuracy of downstream inference by the LLMs. The proof process of Theorems 1 and  2 is provided in Appendix F.1-F.2.